# MomentMix Augmentation with Length-Aware DETR for Temporally Robust Moment Retrieval

### v2

@@ 1 Introduction
In this study, we analyzed the challenges associated with existing methods in short-moment retrieval from both data and model perspectives. From a data perspective, we examined the feature distribution of short moments compared to other moments. As shown in Figure 2, the features of short moments tend to be more concentrated around the mean feature, with 42.9% of samples falling within one standard deviation. In contrast, for other moments, 26.6% of samples fall within one standard deviation of the mean. These observations show that short moments exhibit relatively simple and less diverse feature distributions. On the model side, we analyzed the trends in prediction accuracy by breaking down the model’s final output into center and length components, as illustrated in Figure 3. Although conceptually, moments are defined by (start, end), existing models predict them in the format of (center, length). Interestingly, we found that the accuracy of both center and length predictions for short moments was significantly lower than for other moment types.
Based on the above analysis, we propose a new DETR-based MR framework that addresses the performance degradation when retrieving short moments. Our framework consists of two novel techniques: a data augmentation technique called MomentMix and a Length-Aware Decoder (LAD). Through data analysis, we identified a key limitation in the feature diversity of short moments. It prompts us to design MomentMix, a two-stage mix-based augmentation strategy specifically to enhance the diversity of short-moment samples. In a video sample, we define the temporal moments relevant to the text query as the foreground and the unrelated moments as the background. MomentMix operates in two stages. In the first stage, we create new short foregrounds from rich foreground elements by cutting and mixing. In the next stage, to further diversify these short samples, we utilize portions of other videos as backgrounds, forming varied foreground-background combinations. This increased feature diversity of short moments leads to more robust detection of short moments (see Fig. A1 in Supple.).
@@ 2.1 Moment Retrieval
Figure 2: Data Perspective Analysis. t-SNE [32] visualization of visual features distribution for 50 sampled short moments and 50 non-short moments from QVHighlights train set. Each sampled moment is taken from a distinct video. This plot demonstrates that short moments exhibit fewer and significantly sparser visual features compared to non-short moments, highlighting the limited visual information inherent to their short duration.
@@ 3.1 Motivation
For the data-centric analysis, we examined the statistical characteristics of short moments. As illustrated in QVHighlights [17], while the total number of short moments is comparable to that of other types of moments, the number of videos containing short moments is clearly limited. This led us to hypothesize that short moments might lack diverse contextual representation and could exhibit a narrow distribution in the training data. To test this hypothesis, we compared the feature distribution of short moments to that of other moments using feature visualization. We randomly sampled 50 short moments and 50 non-short moments from the training set and applied t-SNE [32] to visualize the distribution of their visual features. As shown in Figure 2, the distribution of visual features for short moments was concentrated, indicating a significant lack of diversity. This observation suggests that the training data for short moments does not capture a wide range of visual features, leading to suboptimal generalization performance during testing.
@@ 3.2 MomentMix: Augmentation for Short Moment
Second Stage: BackgroundMix. The goal of BackgroundMix is to improve the diversity of the visual background features of augmented short samples from ForegroundMix, thereby strengthening the association between foreground visual features and the text query. To achieve this, we keep the original foreground features while replacing the background with features from various videos. This method provides the model with richer training signals, allowing it to learn various boundaries more effectively for robust short moment retrieval.
@@
This approach ensures that while the backgrounds of the $k$-th sample are augmented with diverse background features, the original foreground remains intact.
@@ 4.3 Ablation Studies and Discussions
Effect of MomentMix. To validate the effectiveness of MomentMix, a novel mixing-based augmentation for moment retrieval (MR) for short moment, we compare it with four baseline augmentations: Gaussian Noise, Random Drop, Random Crop and Random Insert. For Gaussian Noise and Random Drop, we report only the best-performing hyperparameters: adding Gaussian noise with a standard deviation of 0.01 and masking 50% of frame features, respectively. Random Crop and Random Insert were originally proposed for proposal-based long-video MR and adapted by us for proposal-free DETR-based methods. In contrast, our proposed MomentMix uniquely employs foreground mixing (ForegroundMix) and background replacement (BackgroundMix). Specifically, ForegroundMix cuts and mixes foreground segments rather than simply cropping them, while BackgroundMix replaces each background segment independently with clips from different videos, instead of inserting a single foreground segment. As shown in Table 5, MomentMix consistently outperforms baseline augmentations, significantly boosting performance on short moments through enhanced feature diversity (see Fig. A1 in Supple.), while also improving longer moments.
@@
Attention in LAD. To examine whether our LAD attends to the characteristics of each length-specific query as intended, we investigated attention between these queries and encoder features. In Fig. 6, the existing method [25] predicts moments by focusing on the boundary regardless of length, which we refer to as boundary attention of the moment. However, when applying our method, we observed that attention varies according to the characteristics of each length-specific query. For short moments, attention is focused more on the inside of the moment rather than the boundary, which we refer to as inside attention of the moment. On the other side, for very long moments, attention is directed toward the boundary similar to the existing method. This indicates that our query attends to the length rather than the center of the moment appears to be more efficient for detection, hence the boundary attention. Conversely, for short moments, focusing on the center rather than the length is more efficient, resulting in a moment of inside attention.
@@
Figure 7: A qualitative result on QVHighlights. Existing models often fail to accurately distinguish between foreground and background, leading to unsuccessful predictions or missed detections of short moments. In contrast, our model is capable of predicting short moments with greater accuracy and robustness.
@@ 5 Discussion
Future Work. In future work, we plan to explore adapting and extending our augmentation approach to non-DETR architectures. Additionally, conducting further experiments on larger-scale models could provide deeper insights into the scalability and robustness of our method.
@@ Appendix A Additional Ablation Studies
Representation of augmented data. We analyzed the performance degradation of short moments from a data perspective and observed that short-moment visual features exhibit limited diversity. To address this, we applied our MomentMix augmentation strategy to enhance feature diversity. To verify whether MomentMix effectively resolved this issue, we visualized the feature distributions of augmented short moments. As shown in Fig. A1, we found that the previously limited short-moment feature diversity increased significantly, approaching levels comparable to non-short moments. This confirms that MomentMix successfully mitigates the feature diversity issue, leading to improved model performance.
Figure A1: We performed the same analysis as in Fig. 2 on training data with our augmentation. The augmented short moment features lie within the same distribution as the original training features but exhibits greater diversity.
Effect of Length-Aware Decoder. We present the Length-Aware Decoder (LAD), a novel framework designed to improve moment center predictions by conditioning on moment length. Unlike Group-DETR [4] in object dectection, which employs a group-wise one-to-many matching strategy without explicit group definitions, LAD utilizes length class-wise one-to-one matching to generate length-wise expert queries, as shown in Tab. A1. To validate LAD’s effectiveness as a specialized and robust solution for MR, we applied both LAD and Group-DETR to a shared baseline [25].
@@
Table A1: Performance comparison of LAD (Length-Aware Decoder) and Group-DETR on the QVHighlights val set.
Method Short Middle Long Full
R1 mAP R1 mAP R1 mAP R1 mAP
Baseline 4.57 7.77 38.89 43.10 42.62 47.44 41.06 41.00
Group-DETR 4.90 3.84 40.24 40.72 43.14 43.79 42.17 37.97
LAD 8.76 11.01 40.55 45.53 43.69 50.76 43.65 44.48
As shown in Table A1, while Group-DETR improves R1, it suffers a significant drop in mAP. In contrast, LAD achieves substantial gains in both R1 and mAP, effectively addressing performance degradation in short moments through its length-aware mechanisms. These results underscore LAD as a more effective and task-specific approach for MR.
@@ Appendix B Moment Length Class Selection
Method Short Middle Long Full
R1 mAP R1 mAP R1 mAP R1 mAP
Avg. Avg. Avg. @0.5 @0.7 Avg. @0.5 @0.75 Avg.
[…]
Method Short Middle Long Full
R1 mAP R1 mAP R1 mAP R1 mAP
Avg. Avg. Avg. @0.5 @0.7 Avg. @0.5 @0.75 Avg.
[…]
Table A7: Performance when applying a consistent class definition across all datasets. The performance consistently improves.
@@
+ Ours 49.03 50.27 42.01 27.24 59.78 40.48
$^{\textbf{{(+2.26)}}}$ $^{\textbf{{(+4.47)}}}$ $^{\textbf{{(+5.62)}}}$ $^{\textbf{{(+3.92)}}}$ $^{\textbf{{(+0.53)}}}$ $^{\textbf{{(+3.84)}}}$
Effect of consistent class definition. We investigated the performance when using fixed thresholds [10, 30, 70, inf], deviating from the aforementioned approach. As depicted in Tab. A7, this also led to improved performance across all the datasets. This results demonstrates robust performance regardless of class definitions and holds the promise for further enhancement through precise tuning tailored to dataset characteristics.
@@ Appendix D More Qualitative Results
We provide comparisons with other models across a broader range of samples. Through these examples in Figure A4, we can reaffirm that our method exhibits superior accuracy in predicting short moments.
Figure A4: We visualized predictions with confidence scores exceeding 0.7, using an alpha value of 0.5 on QVHighlight val set. ”Ours” refers to UVCOM with our proposed method applied. Existing models frequently struggle to effectively distinguish between foreground and background, resulting in inaccurate predictions or missed detections of short moments. In contrast, our model excels in accurately and robustly predicting short moments.
[…]
[supplement]

### v3

@@ 1 Introduction
In this study, we analyzed the challenges associated with existing methods in short-moment retrieval from both data and model perspectives. From a data perspective, we examined the feature distribution of short moments compared to other moments. As shown in Figure 2, the features of short moments tend to be more substantially concentrated than that of non‑short: the 95% confidence ellipse of the covariance for short moments is markedly smaller. This observation shows that short moments exhibit relatively simple and less diverse feature distributions. On the model side, we analyzed the trends in prediction accuracy by breaking down the model’s final output into center and length components, as illustrated in Figure 3. Although conceptually, moments are defined by (start, end), existing models predict them in the format of (center, length). Interestingly, we found that the accuracy of both center and length predictions for short moments was significantly lower than for other moment types.
Based on the above analysis, we propose a new DETR-based MR framework that addresses the performance degradation when retrieving short moments. Our framework consists of two novel techniques: a data augmentation technique called MomentMix and a Length-Aware Decoder (LAD). Through data analysis, we identified a key limitation in the feature diversity of short moments. It prompts us to design MomentMix, a two-stage mix-based augmentation strategy specifically to enhance the diversity of short-moment samples. In a video sample, we define the temporal moments relevant to the text query as the foreground and the unrelated moments as the background. MomentMix operates in two stages. In the first stage, we create new short foregrounds from rich foreground elements by cutting and mixing. In the next stage, to further diversify these short samples, we utilize portions of other videos as backgrounds, forming varied foreground-background combinations. This increased feature diversity of short moments leads to more robust detection of short moments.
@@ 2.1 Moment Retrieval
Figure 2: Data Perspective Analysis. We analyze QVHighlights val features in ICA [12]. [Left] short moments form a tighter cluster, as indicated by a smaller 95% confidence ellipse of the covariance. [Right] For all Independent Components, the Median Absolute Deviation is greater for non-short moments, as shown by the majority of points lying above $y=x$. Overall, these highlight short moments have the limited visual information.
@@ 3.1 Motivation
For the data-centric analysis, we examined the statistical characteristics of short moments. As illustrated in QVHighlights [18], while the total number of short moments is comparable to that of other types of moments, the number of videos containing short moments is clearly limited. This led us to hypothesize that short moments might lack diverse contextual representation and could exhibit a narrow distribution in the data. To test this hypothesis, we compared the feature distribution of short moments to that of other moments. We analyze visual features in a fixed ICA [12] coordinate system. As shown in Figure 2, on the selected independent components (ICs) the feature distribution of short moments was concentrated, indicating a significant lack of diversity. This pattern generalizes across other components, as all ICs have larger median absolute deviation for non-short moments. These observations suggest that short moments does not capture a wide range of visual features, leading to suboptimal generalization performance.
@@ 3.2 MomentMix: Augmentation for Short Moment
Second Stage: BackgroundMix. The goal of BackgroundMix is to improve the diversity of the visual background features of augmented short samples from ForegroundMix, thereby strengthening the association between foreground visual features and the text query. To achieve this, we keep the original foreground features while replacing the background with features from various videos’ ground truth moment whose queries are semantically dissimilar to the current query. This method provides the model with richer training signals, allowing it to learn various boundaries more effectively for robust short moment retrieval.
@@
To avoid introducing query-related content into the background, we replace backgrounds only with segments from other videos’ ground-truth moments whose queries are semantically dissimilar to the current query.
@@ 4.2 Results
Figure 6: Qualitative results on QVHighlights. Existing models often fail to accurately distinguish between foreground and background, leading to unsuccessful predictions or missed detections of short moments. In contrast, our model is capable of predicting short moments with greater accuracy and robustness.
@@ 4.3 Ablation Studies and Discussions
Effect of MomentMix. To validate the effectiveness of MomentMix, a novel mixing-based augmentation for moment retrieval (MR) for short moment, we compare it with four baseline augmentations: Gaussian Noise, Random Drop, Random Crop and Random Insert. For Gaussian Noise and Random Drop, we report only the best-performing hyperparameters: adding Gaussian noise with a standard deviation of 0.01 and masking 50% of frame features, respectively. Random Crop and Random Insert were originally proposed for proposal-based long-video MR and adapted by us for proposal-free DETR-based methods. In contrast, our proposed MomentMix uniquely employs foreground mixing (ForegroundMix) and background replacement (BackgroundMix). Specifically, ForegroundMix cuts and mixes foreground segments rather than simply cropping them, while BackgroundMix replaces each background segment independently with clips from different videos, instead of inserting a single foreground segment. As shown in Table 5, MomentMix consistently outperforms baseline augmentations, significantly boosting performance on short moments, while also improving longer moments.
@@
Attention in LAD. To examine whether our LAD attends to the characteristics of each length-specific query as intended, we investigated attention between these queries and encoder features. In Fig. 7, the existing method [26] predicts moments by focusing on the boundary regardless of length, which we refer to as boundary attention of the moment. However, when applying our method, we observed that attention varies according to the characteristics of each length-specific query. For the first length class (shorter moments), attention is focused more on the inside of the moment rather than the boundary, which we refer to as inside attention of the moment. On the other side, for the fourth length class (longer moments), attention is directed toward the boundary similar to the existing method. This indicates that our query attends to the length rather than the center of the moment appears to be more efficient for detection, hence the boundary attention. Conversely, for short moments, focusing on the center rather than the length is more efficient, resulting in a moment of inside attention.
@@ 5 Discussion
Future Work. In future work, we plan to explore adapting and extending our augmentation approach to non-DETR architectures. Additionally, to better handle datasets with strong temporal dependencies, we will explore more sophisticated temporal-aware augmentations.
Acknowledgments
This research was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the MSIP (No. RS-2025-00520207 and No. RS-2023-00240135); the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project and RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)); and the IITP grant funded by MSIT and KEIT grant funded by the Korea government (MOTIE) (No. 2022-0-01045, No. 2022-0-00680, and No. RS-2025-02217259, Development of self-evolving AI bias detection-correction-explain platform based on international multidisciplinary governance, 14.3%). We thank Hyogon Ryu and Jihwan Eom for their valuable discussions.
@@ Appendix A Additional Ablation Studies
Figure A1: Pairwise scatter plots for the selected ICs. Diagonals show the 1D density (KDE), while off-diagonals show 2D distributions. The larger 95% covariance ellipses for the Non-short Moment group provide strong evidence of its greater feature diversity.
ICA-Based Feature Diversity Analysis. To investigate the feature distributions of the QVHighlights val set, we project the features into a latent space constructed using PCA and FastICA. To ensure robustness, we filter out unstable components and retain only those showing high consistency across multiple initialization runs. Figure A1 presents the pairwise distributions of the components with the most distinct differences between short and non-short groups. The plot qualitatively demonstrates that non-short moments occupy a broader region in the latent space than short moments, as evidenced by wider kernel density estimates (KDEs) and larger covariance ellipses.
Effect of Length-Aware Decoder. We introduce the Length-Aware Decoder (LAD), a novel framework designed to improve moment center predictions by conditioning on moment length. LAD’s core innovation is a length-wise bipartite matching scheme, which assigns queries to predefined length classes to generate length-expert queries. While LAD builds upon Anchor-DETR, Anchor-DETR’s pattern embeddings alone are not length-aware. Furthermore, LAD’s strategy differs from the length class-wise one-to-one matching of Group-DETR by employing a more precise length class-wise one-to-one matching tailored for moment retrieval, as shown in Fig. A2.
@@
The effectiveness of our approach is validated in Table A1 with two key findings. First, applying LAD’s length-wise matching to the Anchor-DETR baseline yields significant gains in both overall and short-moment mAP, proving the value of explicit length conditioning. Second, in contrast to Group-DETR, which improves R@1 but suffer from significant drop in mAP, LAD achieves substantial improvements in both metrics. These results underscore that LAD’s specialized mechanism effectively addresses performance degradation in challenging short moments, establishing it as a more robust and task-specific solution.
@@
Method Short Middle Long Full
R1 mAP R1 mAP R1 mAP R1 mAP
Baseline 4.57 7.77 38.89 43.10 42.62 47.44 41.06 41.00
Anchor-DETR 4.48 7.43 39.53 43.36 45.4 49.09 42.46 41.55
Group-DETR 4.90 3.84 40.24 40.72 43.14 43.79 42.17 37.97
LAD 8.76 11.01 40.55 45.53 43.69 50.76 43.65 44.48
Table A1: Performance comparison on QVHighlights val set.
@@ Appendix B Moment Length Class Selection
Method Short Middle Long Full
R1 mAP R1 mAP R1 mAP R1 mAP
Avg. Avg. Avg. @0.5 @0.7 Avg. @0.5 @0.75 Avg.
[…]
Method Short Middle Long Full
R1 mAP R1 mAP R1 mAP R1 mAP
Avg. Avg. Avg. @0.5 @0.7 Avg. @0.5 @0.75 Avg.
[…]
@@
Comparison of class-division strategies. We evaluate three substitutes for K-means—equal-interval, equal-count, and Fisher–Jenks variance minimization. As depicted in Tab. A7, every alternative yields substantial gains over the baseline, confirming robustness to how length classes are defined. Across all datasets, K-means consistently achieved the best performance, so we adopt it as the default.
@@
K-means 49.85 50.76 42.21 28.02 61.45 40.22
Equal-Interval 48.86 49.21 41.41 27.99 60.67 40.51
Equal-Count 48.39 50.42 41.54 27.09 58.47 38.66
Fisher-Jenks 47.65 49.58 41.01 26.84 60.19 40.86
Table A7: Performance Comparison of different strategies for moment length class definition. Across all datasets, all strategies improve performance over the baseline; K-means yields the best results overall.
@@ Appendix D More Qualitative Results
We provide comparisons with other models across a broader range of samples. Predictions with confidence scores exceeding 0.7 are visualized with alpha = 0.5 on the QVHighlights val set; “Ours” denotes UVCOM augmented with our method.
As shown in Figure A4, Existing models frequently struggle to effectively distinguish between foreground and background, resulting in inaccurate predictions or missed detections of short moments. In contrast, our model excels in accurately and robustly predicting short moments. our approach yields consistently higher accuracy on short moments.
We also include representative failure cases that are difficult irrespective of length, primarily due to visual–text alignment challenges. Examples are strong window reflections that obscure content and screen-in-screen scenes that induce visual ambiguity as shown in Figure A5.
Figure A4: Qualitative success on short moments with sharper boundaries and fewer spurious hits.
Figure A5: Qualitative failure driven by difficult visual–text alignment.